Culver Company began operations on January 2, 2019. It employs 12 individuals who work 8-hour days and are paid hourly. Each employee earns 11 paid vacation days and 8 paid sick days annually. Vacation days may be taken after January 15 of the year following the year in which they are earned. Sick days may be taken as soon as they are earned; unused sick days accumulate. Additional information is as follows.

Actual Hourly
Wage Rate

Vacation Days Used
by Each Employee

Sick Days Used
by Each Employee

2019

2020

2019

2020

2019

2020

$8 $9 0 10 6 7

Culver Company has chosen to accrue the cost of compensated absences at rates of pay in effect during the period when earned and to accrue sick pay when earned.

Prepare journal entries to record transactions related to compensated absences during 2019 and 2020.

Compute the amounts of any liability for compensated absences that should be reported on the balance sheet at December 31, 2019 and 2020.

2019

2020

Vacation Wages Payable

$enter a dollar amount

$enter a dollar amount

Sick Pay Wages Payable

$enter a dollar amount

$enter a dollar amount
Answer #1

a. Journal entries:

Year Account Debit Credit Note No
2019 Wages Expense $14,592
Vacation Wages Payable $8,448 1
Sick Pay Wages Payable $6,144 2
To accrue expense and liability for compensated absences
2019 Sick Pay Wages Payable $4,608 3
Cash $4,608
To record payment for compensated time when used
2020 Wages Expense $16,416
Vacation Wages Payable $9,504 4
Sick Pay Wages Payable $6,912 5
To accrue expense and liability for compensated absences
2020 Wages Expense $1,152
Vacation Wages Payable $7,680 6
Sick Pay Wages Payable $5,856 7
Cash $14,688 8
To record payment for compensated time when used

Notes:

(1) 12 employees * $8/hr. * 8 hrs./day * 11 days = $8,448

(2) 12 employees * $8/hr. * 8 hrs./day * 8 days = $6,144

(3) 12 employees * $8/hr. * 8 hrs./day * 6 days = $4,608

(4) 12 employees * $9/hr. * 8 hrs./day * 11 days = $9,504

(5) 12 employees * $9/hr. * 8 hrs./day * 8 days = $6,912

(6) 12 employees * $8/hr. * 8 hrs./day * 10 days = $7,680

(7) 12 employees * $8/hr. * 8 hrs./day * ( 8 - 6) days = $1,536

Add: 12 employees * $9/hr. * 8 hrs./day * (7 - 2) days = $4,320

= $1,536 + $4,320 = $5,856

(8) 12 employees * $9/hr. * 8 hrs./day * 10 days = $8,640

Add: 12 employees * $9/hr. * 8 hrs./day * 7 days = $6,048

= $8,640 + $6,048 = $14,688

b. Accrued liability at year-end:

2019 2020
Vacation Wages Payable Sick Pay Wages Payable Vacation Wages Payable Sick Pay Wages Payable
Jan. 1 balance $0 $0 $8,448 $1,536
Add: Accrued $8,448 $6,144 $9,504 $6,912
Less: Paid $0 -$4,608 -$7,680 -$5,856
Dec. 31 balance $8,448 (1) $1,536 (2) $10,272 (3) $2,592 (4)

Notes:

(1) 12 employees * $8/hr. * 8 hrs./day * 11 days = $8,448

(2) 12 employees * $8/hr. * 8 hrs./day * (8 - 6) days = $1,536

(3) 12 employees * $8/hr. * 8 hrs./day * (11 - 10) days = $ 768

Add: 12 employees * $9/hr. * 8 hrs./day * 11 days = $9,504

= $768 + $9,504 = $10,272

(4) 12 employees * $9/hr. * 8 hrs./day * (8 + 8 - 6 - 7) days = $2,592